How do I calculate the resistence of R2 current, which is parallel to R1 (resistence 4 ohm, current 2,25 A). Additional to that I have a resistence R3 which is a series circuit (5 ohm) next to the R1 and R2. They are all on a 24 V battery.


1
Expert's answer
2021-02-11T12:01:11-0500

Current through resistance R3:

Voltage on R1:

"U_1=I_1R_1"

Since R1 and R2 are parallel, then:

"U_1=U_2=I_2R_2"

Total voltage:

"U=U_1+U_3=I_1R_1+I_3R_3"

Then:

"I_3=\frac{U-I_1R_1}{R_3}=\frac{24-2.25\cdot4}{5}=3A"

"I_2=I_3-I_1=3-2.25=0.75A"


"R_2=\frac{I_1R_1}{I_2}=\frac{2.25\cdot4}{0.75}=12" ohm